OpenStudy (anonymous):

Hello I'm looking for help with my algebra...the first question is - the equation 3x+2y=16 describes a line and is written in standard form, (A) solve for the variable y, and (B) use this result to find y with x=-4

OpenStudy (anonymous):

Part A) to solve for variable y is to make y the subject; 3x + 2y = 16 2y =16-3x y =(16-3x)/2 y = 8 - 1.5x Part B) we now substitue x = -4 into the equation in part A ; y = 8 - 1.5(-4) y = 8 + 6 = 14

OpenStudy (anonymous):

in my book it said that the answer for (A) was -3/2x + 8 and (B) was y=14

OpenStudy (anonymous):

that would be right for my solution. y = -1.5x + 8 = -3/2x + 8